# Mode analysis of Nambu-Goldstone modes in U(1) charged first-order relativistic viscous hydrodynamics

### v2

@@ Abstract
We conduct a mode analysis of a general U(1)-charged first-order relativistic hydrodynamics within the framework of the effective field theory of dissipative fluids in a Minkowski background. The most general quadratic action for collective excitations around hydrodynamic solutions — specifically, the Nambu-Goldstone (NG) modes associated with the symmetry-breaking pattern induced by external fields — is derived. It is found that the hydrodynamical frame invariants write the first-order dispersion relations in the low energy limit. In thermodynamical viscous fluids realized by integrating out fast modes, unitarity and local KMS symmetry for its underlying UV theory are guaranteed. Then, we find that first-order hydrodynamics are stable if the null energy condition is satisfied. As the NG modes nonlinearly realize the diffeomorphism symmetries, our mode analysis is valid in any coordinate systems, including Lorentz-transformed references. We also comment on the causal structure for the diffusive modes based on the retarded Green functions for the stochastic NG mode.
arxiv: 1234.56789
@@ 1 Introduction
Relativistic viscous hydrodynamics is a powerful framework for studying many body systems whose fundamental interactions are too complex to solve exactly (see Ref. Kovtun:2012rj for a comprehensive review). In this framework, microscopic details are encapsulated in a few transport coefficients through the derivative expansion. Applications of this approach span a wide range of fields, including cosmology, astrophysics, and ultra-relativistic heavy ion collisions. Although the construction of relativistic hydrodynamics is robust, respecting fundamental symmetries like general covariance and internal symmetries related to charge conservation, there are longstanding issues when including first order corrections, i.e., violation of stability and causality. Stability refers to the behavior of hydrodynamical fluctuations around a thermal state. In the long-time limit, a thermal state should be stationary, and any small perturbations must decay over time. Causality demands that the speed of signal remain bounded by the speed of light. These two properties are essential for any reasonable physical system. Nevertheless, first-order hydrodynamics has been considered to be inconsistent with them. For a more exhaustive summary of the various proposals addressing these issues over the past 80 years, see Ref. Bemfica:2020zjp and references therein.
We study the behavior of perturbations around hydrodynamical solutions in a relativistic framework. To address these problems, it is crucial to clarify the relevant scales in perturbation theory. Realization of a hydrodynamical solution implies that spontaneous symmetry breaking (SSB) of the diffeomorphism symmetries takes place in the presence of a rest frame of fluid. Then such hydrodynamical perturbations are described by the Nambu-Goldstone (NG) modes associated with the diffeomorphism symmetries Crossley:2015evo . From the effective field theory (EFT) point of view, energy density $\rho$ is the energy scale of SSB, $M_{\rm SSB}\sim\rho^{1/4}$, i.e., the scale where the perturbative picture is valid when canonically normalizing the NG mode. $M_{\rm diss.}\sim\eta^{1/3}$, associated with dissipation with shear viscosity $\eta$, typically determined by the mean free path (MFP) in the underlying UV theory. The gradient expansion is an expansion in $\varepsilon=\omega M_{\rm diss.}^{3}/M_{\rm SSB}^{4}$, and first-order hydrodynamics is valid when $\varepsilon\ll 1$. Any unwanted features may arise outside this regime, which implies that it is beyond the applicability range of first-order hydrodynamics Geroch:1995bx . In other words, if theory is subject to unphysical feature in the hydrodynamical regime, it is crucial.
Ref. Hiscock:1985zz studied perturbations in a general hydrodynamical frame, finding that a non-vanishing momentum density $\kappa$ leads to an unstable gapped excitation $e^{(\rho+P)t/\kappa}$, where $\rho$ is the energy density and $P$ is the pressure. This instability is independent of the Fourier wavenumber $k$, implying it manifests on hydrodynamic scales. While eliminating the momentum density in the Landau frame stabilizes the perturbations, the limit $\kappa\to 0$ exacerbates the instability. Additionally, they found unstable modes even for $\kappa=0$ in non-comoving frames.
In this paper, we claim that these instabilities stem from an inconsistent treatment of the gradient expansion. Proper order counting in $\varepsilon$ eliminates these unstable modes in the hydrodynamic regime. Thus, first-order hydrodynamics does not have inherent instability as long as the system remains within the hydrodynamic regime. This fact might be widely acknowledged. For example, Ref. Kovtun:2012rj explains that the Green functions must be expanded within the validity regime of first-order hydrodynamics (See the last paragraph in their section 2.). Nevertheless, misunderstandings about hydrodynamic instability may persist due to unclear presentations of the correct gradient expansion.
A common error involves solving dispersion relations with higher-order derivative corrections before taking the low-energy limit. This results in multiple solutions (e.g., $n$ solutions for an $n$-th order equation of motion). However, higher-order derivative corrections must be considered as corrections to the lowest order derivative theory, or free theory i.e., perturbation theory around perfect fluid. Hence, higher order derivative terms should be truncated before solving for the roots, as $\mathcal{O}(\varepsilon^{2})$ terms in the equations of motion are unreliable when the constitutive relation is truncated at $\mathcal{O}(\varepsilon)$. Then, solutions are found in perturbative expansion in $\varepsilon$. Similar caveats apply to other effective theories, as noted in, for instance, the footnote of Ref. Weinberg:2008hq .
Recent studies have proposed that specific hydrodynamical frames can address these issues hydrodynamical frames that differ from both the traditional Landau and Eckart frames Bemfica:2017wps ; Kovtun:2019hdm ; Bemfica:2019knx . However, the freedom in hydrodynamical frame choice is nothing but the ambiguity in the field redefinition of hydrodynamical solutions in the generating functional, which should not be physical in the usual sense. From this viewpoint, we adopt hydrodynamical frame independence as a guiding principle for our mode analysis since stability and causality are fundamental physical properties. Hence, this work will take a different approach.
In this paper, we resolve the instability issue within a generic U(1) charged fluid framework, focusing on hydrodynamical frame-independence. We also address the instability caused by Lorentz boost for $\kappa=0$ Hiscock:1985zz , which is less commonly discussed in the literature (see recent discussions in Refs. Bemfica:2020zjp ; Gavassino:2023myj ). Using the EFT of dissipative fluids Crossley:2015evo ; Glorioso:2017fpd ; Liu:2018kfw , we write the action of hydrodynamic modes as the nonlinear realization of general coordinate transformations and global U(1) phase transformation. This formalism resolves the issue of Lorentz boost by construction. We aim to show that, as long as the system stays within the hydrodynamic regime, there are no inherent instabilities in first-order hydrodynamics, consistent with the claims in Ref. Geroch:1995bx .
Causality is another issue in first-order hydrodynamics. Though the theory is relativistic, the sound speed depends solely on thermodynamic variables and is not necessarily constrained to be subluminal. However, the possibility of superluminal propagation raises concerns. Recently, Ref. Heller:2022ejw ; Gavassino:2023myj demonstrated that stable modes in general quantum field theories are always subluminal, based on the analytic structure of retarded Green’s functions. Their conclusions directly apply to our setup since the action of hydrodynamic modes, i.e., Nambu-Goldstone modes, is not different from those in standard quantum field theories. We will also comment on the causality issue of the front velocity of the diffusive modes from this perspective.
@@
We organize this paper as follows. First, we construct the generating functional for a global U(1) theory by formally integrating out the fast modes. The form of this functional is constrained by various conditions inherent in the UV theory, such as unitarity and the Kubo-Martin-Schwinger (KMS) symmetry, as discussed in Sec. 2. In this section, we also review the ambiguity related to field redefinitions and rest frames.
In Sec. 3, we introduce collective modes, i.e., the hydrodynamical modes, into the UV generating functional as the nonlinear realization of symmetries associated with external fields. We derive the quadratic action for the NG modes and eliminate redundant variables by solving the constraint equations, such as number conservation and the continuity equation.
@@
Sec. 4 is devoted to the mode analysis for neutral and U(1)-charged fluids. We carefully determine the dispersion relation while keeping track of the order in the gradient expansion. In Sec. 5, we compute the Green’s functions for the most general U(1)-charged fluid. By focusing on the analytical structure of the retarded Green’s function, we derive the bound on the sound speed and discuss the causality issue.
Finally, we present our conclusions in the last section. Several useful equations are summarized in App. A, and a review of one-dimensional Brownian motion in a harmonic potential is provided in App. B.
@@ 2 Realization of hydrodynamical solutions
In this paper, we are interested in the dynamics of a hydrodynamical mode, which is a perturbation around a hydrodynamical solution. In this section, we begin by discussing the realization of a hydrodynamical solution within a microscopic theory underlying hydrodynamics. In the next section, the collective modes will be introduced as the NG modes associated with the spontaneous symmetry breaking around hydrodynamical solutions.
@@ 2.1 Generating functional of UV model
Let us first consider a UV theory of a complex dynamical variable with a global U(1) symmetry. A hydrodynamical system arises when taking the average of this microscopic system. Hereafter, we call the UV variable the “fast mode”. Couple the fast mode to the spacetime metric $g$ and external vector field $A$. Then, the variation of the UV action $S$ with respect to the external fields introduces the microscopic energy-momentum tensor and the conserved current:
$\hat{T}^{\mu\nu}\equiv\frac{2}{\sqrt{-\det g}}\frac{\delta S}{\delta g_{\mu\nu}},~{}\hat{J}^{\mu}\equiv\frac{1}{\sqrt{-\det g}}\frac{\delta S}{\delta A_{\mu}}.$ (1)
Then, the correlation functions of the energy-momentum tensor and the conserved current are given as
$T^{\mu\nu}$ $=\langle\hat{T}^{\mu\nu}\rangle,$ (2)
$J^{\mu}$ $=\langle\hat{J}^{\mu}\rangle,$ (3)
$W^{\mu\nu\rho\sigma}(x,y)$ $=\langle\hat{T}^{\mu\nu}(x)\hat{T}^{\rho\sigma}(y)\rangle,$ (4)
$X^{\mu\nu\rho}(x,y)$ $=\langle\hat{T}^{\mu\nu}(x)\hat{J}^{\rho}(y)\rangle,$ (5)
$Y^{\mu\rho}(x,y)$ $=\langle\hat{J}^{\mu}(x)\hat{J}^{\rho}(y)\rangle.$ (6)
These averages mean the path integral of the fast mode defined from $S$ in the Schwinger-Keldysh formalism with an initial state. At this stage, $T,J,\cdots$ are macroscopic quantities obtained by averaging the microscopic energy momentum tensor and current with respect to the path integral for a given initial condition. Let us assume the two-point or higher-order correlation functions are approximately written as
$W^{\mu\nu\rho\sigma}(x,y)=W^{\mu\nu\rho\sigma}(x)\frac{\delta(x-y)}{\sqrt{-\det g}},$ (7)
and $X$ and $Y$ are also written in a similar way. The non local contributions will be described by long-lived excitations around a hydrodynamical solution discussed in the next section. Given a list of correlation functions, we may reconstruct the generating functional $e^{iI}$ as
@@
$+\frac{i}{4}\bar{g}_{\mu\nu}W^{\mu\nu\rho\sigma}\bar{g}_{\rho\sigma}+i\bar{g}_{\mu\nu}X^{\mu\nu\rho}\bar{A}_{\rho}+i\bar{A}_{\mu}Y^{\mu\rho}\bar{A}_{\rho}+\cdots\Bigg{]}.$ (8)
Note that the over-bared external fields are the noise counterparts in the Keldysh basis in the Schwinger-Keldysh formalism. The physical external fields are contained in $T,J,\cdots$. The generating functional is local due to the assumption (7). Note that the correlation functions are determined by the path integral, so $T,J,\cdots$ are hydrodynamical solutions.
The actual integration is complex in general. However, one can directly write the correlation functions based on symmetry structure with undetermined coefficients, which is called constitutive relation. Let us write the correlation functions for a general U(1) charged fluid. A hydrodynamical solution is realized in the presence of a time-like 4-vector $\beta^{\mu}$, and external fields $g_{\mu\nu}$ and $A_{\mu}$. As $\beta^{\mu}$ splits external fields into 3+1 form, we can find a convenient set of variables as follows. The magnitude of $\beta^{\mu}$ is identified with the local temperature $\beta\equiv\sqrt{-g_{\mu\nu}\beta^{\mu}\beta^{\nu}}$, and then the fluid 4-velocity $u^{\mu}\equiv\beta^{\mu}/\beta$. The temporal component of the vector field is identified with the chemical potential $\mu\equiv u^{\mu}A_{\mu}$. Finally, the external fields perpendicular to $u^{\mu}$ are $\gamma_{\mu\nu}\equiv u_{\mu}u_{\nu}+g_{\mu\nu}$, and $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\mu}A_{\nu}$. We write $\Psi=\{T,J,W,X,Y\}$ in terms of $\Phi=\{\beta^{\mu},g_{\mu\nu},A_{\mu}\}=\{\beta,u^{\mu},\mu,F_{\mu\nu},\gamma_{\mu\nu}\}$. We consider the most general forms, order by order, in the covariant derivative:
$\Psi=\sum_{n=0}\Psi_{(n)},$ (9)
where subscript $n$ counts the number of covariant derivatives. Note that $\nabla$ is always projected to $u^{\mu}$ or $\gamma^{\mu\nu}$. $F$ is the first order in $\nabla$, and the rest of the variables in $\Phi$ are the zeroth order.
@@
In the lowest order, the spacetime indexes are only represented by $u^{\mu}$ and $\gamma^{\mu\nu}$. The most general $\Psi_{(0)}$ are
$T^{\mu\nu}_{(0)}$ $=\rho u^{\mu}u^{\nu}+P\gamma^{\mu\nu},$ (10)
$J^{\mu}_{(0)}$ $=nu^{\mu},$ (11)
$\begin{split}W_{(0)}^{\mu\nu\rho\sigma}&=w_{1}u^{\mu}u^{\nu}u^{\rho}u^{\sigma}+w_{2}(\gamma^{\mu\nu}u^{\rho}u^{\sigma}+\gamma^{\rho\sigma}u^{\mu}u^{\nu})\\ &+w_{3}(\gamma^{\mu\rho}u^{\nu}u^{\sigma}+\gamma^{\nu\rho}u^{\mu}u^{\sigma}+\gamma^{\mu\sigma}u^{\nu}u^{\rho}+\gamma^{\nu\sigma}u^{\mu}u^{\rho})\\ &+w_{4}\gamma^{\mu\nu}\gamma^{\rho\sigma}+w_{5}\left(\gamma^{\mu\rho}\gamma^{\nu\sigma}+\gamma^{\nu\rho}\gamma^{\mu\sigma}-\frac{2}{3}\gamma^{\mu\nu}\gamma^{\rho\sigma}\right).\end{split}$ (12)
$\begin{split}X_{(0)}^{\mu\nu\rho}&=\beta x_{1}u^{\mu}u^{\nu}u^{\rho}+\beta x_{2}\gamma^{\mu\nu}u^{\rho}+\beta x_{3}(\gamma^{\mu\rho}u^{\nu}+\gamma^{\nu\rho}u^{\mu}).\end{split}$ (13)
$\begin{split}Y_{(0)}^{\mu\rho}&=\beta^{2}y_{1}u^{\mu}u^{\rho}+\beta^{2}y_{2}\gamma^{\mu\rho}.\end{split}$ (14)
Note that we assumed Eq. (7). The index symmetry in correlation functions imply $W^{\mu\nu\rho\sigma}=W^{\nu\mu\rho\sigma}=W^{\rho\sigma\mu\nu}$, $X^{\mu\nu\rho}=X^{\nu\mu\rho}$ and $Y^{\mu\nu}=Y^{\nu\mu}$. The expansion coefficients $\rho$, $P$, $n$ $w$, $x$ and $y$ are functions of $\beta$ and $\mu$, which are undetermined within the effective theory. The mass dimensions of the coefficients are given as $[T]=M^{4}$, $[J]=M^{3}$ $[W]=M^{4}$, $[X]=M^{3}$ and $[Y]=M^{2}$.
@@
$\begin{split}T^{\mu\nu}_{(1)}&=\rho_{(1)}u^{\mu}u^{\nu}+P_{(1)}\gamma^{\mu\nu}+q_{(1)\sigma}(u^{\mu}\gamma^{\nu\sigma}+u^{\nu}\gamma^{\mu\sigma})-\Sigma_{(1)}^{\mu\nu}\end{split},$ (15)
$\begin{split}J^{\mu}_{(1)}&=n_{(1)}u^{\mu}+\gamma^{\mu\nu}j_{(1)\nu},\end{split}$ (16)
where we defined the following first-order coefficients:
$\rho_{(1)}$ $=\beta^{-1}\epsilon u^{\rho}\nabla_{\rho}\beta-\lambda_{1}\nabla_{\rho}u^{\rho}-\nu_{4}u^{\rho}\nabla_{\rho}(\beta\mu),$ (17)
$P_{(1)}$ $=\beta^{-1}{\lambda_{2}}u^{\rho}\nabla_{\rho}\beta-\zeta\nabla_{\rho}u^{\rho}-\nu_{5}u^{\rho}\nabla_{\rho}(\beta\mu),$ (18)
$n_{(1)}$ $=\nu_{1}u^{\rho}\nabla_{\rho}\beta-\beta\nu_{2}\nabla_{\rho}u^{\rho}-\beta\nu_{3}u^{\rho}\nabla_{\rho}(\beta\mu),$ (19)
$q_{(1)\mu}$ $=\beta^{-1}\kappa_{1}\nabla_{\mu}\beta-\kappa_{2}u^{\rho}\nabla_{\rho}u_{\mu}-\tau_{3}\nabla_{\mu}(\beta\mu)-\beta\tau_{4}u^{\rho}F_{\rho\mu},$ (20)
$j_{(1)\mu}$ $=\tau_{1}\nabla_{\mu}\beta-\beta\tau_{2}u^{\rho}\nabla_{\rho}u_{\mu}-\beta\chi_{1}\nabla_{\rho}(\beta\mu)-\beta^{2}\chi_{2}u^{\rho}F_{\rho\mu},$ (21)
@@
$\Sigma_{(1)}^{\mu\nu}$ $=\eta\gamma^{\mu\alpha}\gamma^{\nu\beta}\left(\nabla_{\beta}u_{\alpha}+\nabla_{\alpha}u_{\beta}-\frac{2}{3}\gamma_{\alpha\beta}\nabla_{\gamma}u^{\gamma}\right).$ (22)

### v3

@@ Abstract
We conduct a mode analysis of a general $U(1)$-charged first-order relativistic hydrodynamics within the framework of effective field theory for dissipative fluids in flat Minkowski spacetime. We derive the most general quadratic action for hydrodynamic modes, including stochastic noise, and analyze the corresponding dispersion relations in a consistent gradient expansion. We argue that spontaneous breaking of spacetime symmetry arises in the presence of a local thermal state specified by a local timelike four-vector. We demonstrate that hydrodynamical perturbations can be identified as Nambu–Goldstone (NG) modes, analogous to their embedding in global $U(1)$-invariant theories. We find that frame-invariant combinations of hydrodynamic transport coefficients determine the first-order dispersion relations in the low-energy limit, making the mode analysis manifestly independent of the choice of hydrodynamic frame. Assuming local Kubo–Martin–Schwinger (KMS) symmetry and unitarity of the underlying UV theory, we show that first-order hydrodynamics is stable if the enthalpy density is positive.
@@ 1 Introduction
Relativistic viscous hydrodynamics is a widely used framework for describing many-body systems whose microscopic interactions are too complex to solve exactly (see Ref. [1] for a comprehensive review). In this approach, detailed microscopic information is encapsulated in a small number of transport coefficients organized via a derivative expansion. Applications of relativistic hydrodynamics range from cosmology and astrophysics to the physics of ultra-relativistic heavy-ion collisions [2, 3, 4, 5, 6, 7, 8].
While the formulation of relativistic hydrodynamics is grounded in general principles—such as diffeomorphism invariance and internal symmetries like charge conservation—the inclusion of dissipative effects raises important questions regarding stability and causality [9, 10]. In particular, stability refers to the requirement that perturbations around thermal equilibrium should decay over time, while causality demands that signal propagation be bounded by the speed of light. Both are fundamental requirements for a physically consistent effective theory.
It has been commonly argued that first-order hydrodynamics may violate these conditions. For instance, Ref. [11] shows that while the Landau frame admits stable hydrodynamic modes in a comoving inertial frame, small deviations from this specific frame can lead to exponential instabilities. Such observations have motivated the development of extended frameworks such as the Müller–Israel–Stewart (MIS) theory [12, 13], which aim to restore stability and causality by incorporating second-order corrections. These frameworks introduce additional transport coefficients and dynamical variables, but also come with increased complexity [14]. However, since different hydrodynamical frame choices in the first order theories correspond to field redefinitions in the thermodynamic variables, physical observables should remain invariant under such transformations. When instabilities arise in specific frames, this may indicate a misapplication of the analysis rather than a fundamental inconsistency in the theory itself.
To investigate this point, we study perturbations around hydrodynamic backgrounds within a relativistic EFT framework [15]. A key element of this approach is the identification of dynamical fluctuations as Nambu–Goldstone (NG) modes. Realizing a hydrodynamic background can be interpreted, in a generalized sense, as spontaneous symmetry breaking (SSB): the presence of a fluid rest frame selects a preferred coordinate system. In analogy with phonons in solids, which are NG modes of broken spatial translation symmetry, collective excitations in fluids may also be interpreted as NG modes [16, 17] (See also Ref. [18] for a recent proposal about the spontaneous symmetry breaking in hydrodynamics.). This symmetry-based approach has been successfully applied to construct effective actions with manifest covariance, providing a systematic framework to incorporate the noise sector and to impose fluctuation-dissipation relations through the KMS condition at the action level [19, 20]. These features make the mode analysis particularly transparent and robust.
We show that some commonly reported instabilities can be traced to mismatches in order counting between the constitutive relations and the perturbative treatment of dynamical modes when solving the constraint equations for the perturbation variables. Similar viewpoints have been expressed in the literature; for example, Ref. [1] notes that Green’s functions should be used only within the regime of validity of first-order hydrodynamics (see the end of their section 2). Recent works have also re-examined the consistency and well-posedness of first-order theories in a different perspective [21, 22, 23]. Here, we aim to clarify this point by performing an explicit analysis with consistent order counting of derivatives in a manifestly covariant framework without hydrodynamical frame dependence, thereby providing a more unified perspective on the stability of relativistic hydrodynamics.
@@
This paper is organized as follows. In Sec. 2, we construct the generating functional for a global $U(1)$ theory underlying hydrodynamics by formally integrating out fast modes. The structure of this functional is constrained by unitarity, KMS symmetry, and other properties of the UV theory. In Sec. 3, we also review the role of field redefinitions and rest-frame ambiguities.
In Sec. 4, we examine the symmetry breaking structure in the presence of a local thermal ensemble. We introduce hydrodynamical modes as the NG modes associated with the spacetime and $U(1)$ symmetries. We derive their quadratic action and eliminate redundant degrees of freedom using the conservation laws.
@@
Section 5 is devoted to mode analysis in both neutral and $U(1)$-charged fluids. We carefully derive the dispersion relations while keeping track of the gradient expansion. In Sec. 6, we compute Green’s functions for the general $U(1)$-charged fluid and examine the analytic structure of the retarded propagators, deriving sound speed bounds and causality constraints.
We conclude with a summary and outlook. For the reader’s convenience, we provide a translation table comparing our hydrodynamical notation with that of Ref. [23] in App. A. Appendix B collects useful thermodynamic identities, and App. C provides a pedagogical review of Brownian motion in a harmonic potential.
@@ 2 Realization of hydrodynamical solutions
We adopt the EFT formulation of dissipative hydrodynamics developed in Ref. [15], which systematically incorporates microscopic constraints such as unitarity and the Kubo-Martin-Schwinger (KMS) condition. While this framework reproduces conventional perturbative results at the linear level (see, e.g., Ref. [23]), it offers several additional advantages:
  • The KMS condition and unitarity are imposed directly on the generating functional, providing a more fundamental alternative to the local second law of thermodynamics. These constraints apply at the level of the effective action itself. When considering a general hydrodynamical frame, the action formalism makes the relationship between the coefficients explicit.
  • Constraint equations for perturbation variables can be imposed consistently up to the order of noise contributions when derived from an action, as opposed to classical equations of motion.
  • General EFT principles apply: the perturbation variables are understood as nonlinear realizations of spontaneously broken symmetries. This renders the treatment of boost transformations trivial and ensures manifest covariance throughout the analysis.
  • The extension to nonlinear interactions in future work is systematic and straightforward within this formalism.
In this section, we demonstrate how a hydrodynamical solution arises from a microscopic theory, providing the groundwork for constructing an EFT of fluctuations around such backgrounds.
This section serves two purposes: to clarify the internal consistency of our approach by comparing it with existing hydrodynamical frameworks, and to establish the notational and conceptual groundwork for the effective action developed later in this paper. Readers primarily interested in the construction and analysis of the hydrodynamical modes may safely skip to Section 4.
@@ 2.1 Generating functional of UV model
Let us first consider a UV theory of a complex dynamical variable with a global U(1) symmetry. A hydrodynamical system can be obtained by coarse-graining over this microscopic system. Hereafter, we refer to the microscopic degrees of freedom as “fast modes”. We couple the fast modes to the spacetime metric $g$ and an external vector field $A$. The variation of the UV action $S$ with respect to these external fields defines the microscopic energy-momentum tensor and conserved current as
$\hat{T}^{\mu\nu}\equiv\frac{2}{\sqrt{-\det g}}\frac{\delta S}{\delta g_{\mu\nu}},~\hat{J}^{\mu}\equiv\frac{1}{\sqrt{-\det g}}\frac{\delta S}{\delta A_{\mu}}.$ (2.1)
The corresponding macroscopic observables are then given by
$T^{\mu\nu}$ $=\langle\hat{T}^{\mu\nu}\rangle,$ (2.2)
$J^{\mu}$ $=\langle\hat{J}^{\mu}\rangle,$ (2.3)
$W^{\mu\nu\rho\sigma}(x,y)$ $=\frac{1}{2}\langle\hat{T}^{\mu\nu}(x)\hat{T}^{\rho\sigma}(y)\rangle,$ (2.4)
$X^{\mu\nu\rho}(x,y)$ $=\frac{1}{2}\langle\hat{T}^{\mu\nu}(x)\hat{J}^{\rho}(y)\rangle,$ (2.5)
$Y^{\mu\rho}(x,y)$ $=\frac{1}{2}\langle\hat{J}^{\mu}(x)\hat{J}^{\rho}(y)\rangle.$ (2.6)
These expectation values are defined through the path integral over fast modes using the Schwinger-Keldysh formalism with an appropriate initial state. At this stage, $T$, $J$, etc., on the left-hand side represent the macroscopic observables obtained by averaging over the microscopic energy-momentum tensor and current under a given initial condition.
We assume that the two-point and higher-order correlation functions are approximately local in space-time:
$W^{\mu\nu\rho\sigma}(x,y)=W^{\mu\nu\rho\sigma}(x)\frac{\delta(x-y)}{\sqrt{-\det g}},$ (2.7)
and similarly for $X$ and $Y$. Nonlocal correlations will be encoded in long-lived excitations, which will be discussed in the next section.
Given this set of correlation functions, one can reconstruct the generating functional $e^{iI}$ as
@@
$+\frac{i}{4}\bar{g}_{\mu\nu}W^{\mu\nu\rho\sigma}\bar{g}_{\rho\sigma}+i\bar{g}_{\mu\nu}X^{\mu\nu\rho}\bar{A}_{\rho}+i\bar{A}_{\mu}Y^{\mu\rho}\bar{A}_{\rho}+\cdots\Bigg{]}.$ (2.8)
Here, the barred external fields $\bar{g}_{\mu\nu}$ and $\bar{A}_{\mu}$ represent auxiliary fields in the Keldysh basis of the Schwinger-Keldysh formalism. [^1] We avoid using the conventional subscript $A$ to prevent confusion with the vector field $A$ and spacetime indices. The unbarred external fields couple to physical observables such as $T^{\mu\nu}$ and $J^{\mu}$. Due to the assumption (2.7), the generating functional is local in spacetime. The correlation functions $T$, $J$, etc. are determined by the path integral and characterize the hydrodynamical state of the system.
The actual integration is complex in general. However, one can write the correlation functions based on symmetry structure with undetermined coefficients, which are called constitutive relations. Let us write the correlation functions for a general U(1)-charged fluid. A hydrodynamic solution is realized in the presence of a timelike 4-vector $\beta^{\mu}$, and external fields $g_{\mu\nu}$ and $A_{\mu}$. As $\beta^{\mu}$ splits the external fields into 3+1 form, we can find a convenient set of variables as follows. The temporal component of the metric is identified with the local temperature $\beta\equiv\sqrt{-g_{\mu\nu}\beta^{\mu}\beta^{\nu}}$. We normalize $\beta^{\mu}$ by $\beta$ and define the fluid 4-velocity $u^{\mu}\equiv\beta^{\mu}/\beta$. The temporal component of the vector field divided by $\beta$ is identified with the chemical potential $\mu\equiv u^{\mu}A_{\mu}$. Finally, the external fields perpendicular to $u^{\mu}$ are $\gamma_{\mu\nu}\equiv g_{\mu\nu}+u_{\mu}u_{\nu}$, and $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$. We write $\Psi=\{T,J,W,X,Y\}$ in terms of $\Phi=\{\beta^{\mu},g_{\mu\nu},A_{\mu}\}=\{\beta,u^{\mu},\mu,F_{\mu\nu},\gamma_{\mu\nu}\}$. We consider the most general forms, order by order, in the covariant derivative:
$\Psi=\sum_{n=0}\Psi_{(n)},$ (2.9)
where the subscript $n$ counts the number of covariant derivatives. Note that $\nabla$ is always projected to $u^{\mu}$ or $\gamma^{\mu\nu}$. $F$ is of first order in $\nabla$, and the rest of the variables in $\Phi$ are of zeroth order.
@@
In the lowest order, the spacetime indices are only represented by $u^{\mu}$ and $\gamma^{\mu\nu}$. The most general $\Psi_{(0)}$ are
$T^{\mu\nu}_{(0)}$ $=\rho u^{\mu}u^{\nu}+P\gamma^{\mu\nu},$ (2.10)
$J^{\mu}_{(0)}$ $=nu^{\mu},$ (2.11)
$\begin{split}W_{(0)}^{\mu\nu\rho\sigma}&=w_{1}u^{\mu}u^{\nu}u^{\rho}u^{\sigma}+w_{2}(\gamma^{\mu\nu}u^{\rho}u^{\sigma}+\gamma^{\rho\sigma}u^{\mu}u^{\nu})\\ &+w_{3}(\gamma^{\mu\rho}u^{\nu}u^{\sigma}+\gamma^{\nu\rho}u^{\mu}u^{\sigma}+\gamma^{\mu\sigma}u^{\nu}u^{\rho}+\gamma^{\nu\sigma}u^{\mu}u^{\rho})\\ &+w_{4}\gamma^{\mu\nu}\gamma^{\rho\sigma}+w_{5}\left(\gamma^{\mu\rho}\gamma^{\nu\sigma}+\gamma^{\nu\rho}\gamma^{\mu\sigma}-\frac{2}{3}\gamma^{\mu\nu}\gamma^{\rho\sigma}\right).\end{split}$ (2.12)
$\begin{split}X_{(0)}^{\mu\nu\rho}&=\beta x_{1}u^{\mu}u^{\nu}u^{\rho}+\beta x_{2}\gamma^{\mu\nu}u^{\rho}+\beta x_{3}(\gamma^{\mu\rho}u^{\nu}+\gamma^{\nu\rho}u^{\mu}).\end{split}$ (2.13)
$\begin{split}Y_{(0)}^{\mu\rho}&=\beta^{2}y_{1}u^{\mu}u^{\rho}+\beta^{2}y_{2}\gamma^{\mu\rho}.\end{split}$ (2.14)
Note that we assumed Eq. (2.7). The index symmetry in correlation functions implies $W^{\mu\nu\rho\sigma}=W^{\nu\mu\rho\sigma}=W^{\rho\sigma\mu\nu}$, $X^{\mu\nu\rho}=X^{\nu\mu\rho}$, and $Y^{\mu\nu}=Y^{\nu\mu}$. The expansion coefficients $\rho$, $P$, $n$, $w$, $x$ and $y$ are functions of $\beta$ and $\mu$, which are undetermined within the effective theory. The mass dimensions of the coefficients are given as $[T]=M^{4}$, $[J]=M^{3}$, $[W]=M^{4}$, $[X]=M^{3}$, and $[Y]=M^{2}$.
@@
$\begin{split}T^{\mu\nu}_{(1)}&=\rho_{(1)}u^{\mu}u^{\nu}+P_{(1)}\gamma^{\mu\nu}+q_{(1)\sigma}(u^{\mu}\gamma^{\nu\sigma}+u^{\nu}\gamma^{\mu\sigma})-\Sigma_{(1)}^{\mu\nu}\end{split},$ (2.15)
$\begin{split}J^{\mu}_{(1)}&=n_{(1)}u^{\mu}+\gamma^{\mu\nu}j_{(1)\nu},\end{split}$ (2.16)
where we define the following first-order coefficients:
$\rho_{(1)}$ $=\beta^{-1}\epsilon u^{\rho}\nabla_{\rho}\beta-\lambda_{1}\nabla_{\rho}u^{\rho}-\nu_{4}u^{\rho}\nabla_{\rho}(\beta\mu),$ (2.17)
$P_{(1)}$ $=\beta^{-1}{\lambda_{2}}u^{\rho}\nabla_{\rho}\beta-\zeta\nabla_{\rho}u^{\rho}-\nu_{5}u^{\rho}\nabla_{\rho}(\beta\mu),$ (2.18)
$n_{(1)}$ $=\nu_{1}u^{\rho}\nabla_{\rho}\beta-\beta\nu_{2}\nabla_{\rho}u^{\rho}-\beta\nu_{3}u^{\rho}\nabla_{\rho}(\beta\mu),$ (2.19)
$q_{(1)\mu}$ $=\beta^{-1}\kappa_{1}\nabla_{\mu}\beta-\kappa_{2}u^{\rho}\nabla_{\rho}u_{\mu}-\tau_{3}\nabla_{\mu}(\beta\mu)-\beta\tau_{4}u^{\rho}F_{\rho\mu},$ (2.20)
$j_{(1)\mu}$ $=\tau_{1}\nabla_{\mu}\beta-\beta\tau_{2}u^{\rho}\nabla_{\rho}u_{\mu}-\beta\chi_{1}\nabla_{\rho}(\beta\mu)-\beta^{2}\chi_{2}u^{\rho}F_{\rho\mu},$ (2.21)
@@
$\Sigma_{(1)}^{\mu\nu}$ $=\eta\gamma^{\mu\alpha}\gamma^{\nu\beta}\left(\nabla_{\beta}u_{\alpha}+\nabla_{\alpha}u_{\beta}-\frac{2}{3}\gamma_{\alpha\beta}\nabla_{\gamma}u^{\gamma}\right).$ (2.22)